OpenStudy (anonymous):

Write the equation 3x+4y-7=0 in normal form. I'll mark the answer I got but it doesn't look right to me. 3/5x+4/5y+7/5=0 -3/5x-4/5y-7/5=0 -3/5x-4/5y+7/5=0 3/5x+4/5y-7/5=0<-- Please explain what normal form is because I'm confused.

OpenStudy (dumbcow):

haha so am i.....@Loser66 what is normal form?

OpenStudy (dumbcow):

apparently it involves dividing by 5 ...

OpenStudy (loser66):

I become dummy with this question, :)

OpenStudy (anonymous):

Yeah I don't know what to do here, I tried looking up normal form but I cant find anything that explains it....we're all stumped thats great....maybe my teacher miswrote the question. XD

hartnn (hartnn):

most probably, normal = standard.

OpenStudy (anonymous):

but isn't the original equation in standard?

hartnn (hartnn):

yes, it is

OpenStudy (loser66):

to me, Yes

hartnn (hartnn):

why only 5 is divided ?? \(\sqrt{3^3+4^2} =5\)

OpenStudy (dumbcow):

yes it is ... the only thing i would say is that if you make the line a vector then dividing by 5 makes it a unit vector

hartnn (hartnn):

the answer you choose is the best one. other choices can never be answer to this

OpenStudy (dumbcow):

i think your answer is right, dont change any of the signs

OpenStudy (anonymous):

Hartnn I don't know, this is all the question states. Dumbcow, thats a good idea I'll see what I can do....although Im terrible with vectors XD

OpenStudy (anonymous):

Okay, thats what I thought because changing the signs would mess up the original equation.

hartnn (hartnn):

yes, good thinking :)

OpenStudy (anonymous):

Oh I found the definition! Normal Form is x cos theta+y sin theta- p=0 Ugh more trig functions, I'm not even taking trig.

OpenStudy (anonymous):

Ax+Bx+C=0 can be changed to Normal form by dividing the terms by +/-sqrtA^2+B^2 with the sign chosen to oppose the sign of C. Sooo...in this case... 3^3+4^2=25 sqrt25=5 3/5x+4/5y-7/5=0 So I was right I think...

hartnn (hartnn):

nice!

OpenStudy (anonymous):

Thanks for the help! and now you guys know too so you can help someone else now!

hartnn (hartnn):

Exactly! thats the point of collaborative studying! :) the very purpose of OpenStudy, learn and help :)